OpenStudy (anonymous):

what is the integration of 1/(x+2)(sqr of x+3)

OpenStudy (anonymous):

whats the solution for this? which method to use?

myininaya (myininaya):

do u=x+2

myininaya (myininaya):

u+1=x+3

myininaya (myininaya):

du=dx

myininaya (myininaya):

int(sqrt(u+1)/sqrt(u^2),u)=int(u^(-1/2)+1/u, u)

myininaya (myininaya):

you did mean for that sqrt(x+3) to be in the numerator correct?

OpenStudy (anonymous):

I thought he meant denominator

myininaya (myininaya):

I wasn't sure since there wasn't brackets around the whole denominator symbolizing it is the denominator

myininaya (myininaya):

If he then mean it to be in the denominator, he could use partial fractions after that substitution above

myininaya (myininaya):

thinker, have you done partial fractions in calculus?

myininaya (myininaya):

oh hes not here

myininaya (myininaya):

see you late krb

myininaya (myininaya):

later*

OpenStudy (anonymous):

goodnite!

OpenStudy (anonymous):

the squareroot part in in the denominator itself..